Your professor created a standard curve with a best-fit line. The equation of the line is y=0.245x+0.342.

a) Calculate the concentration (in %) for the following absorbance values.

0.342, 3.861, 1.462

b) What absorbance would you expect to measure for the following concentrations

18%, 4%, 50%

c) Draw a graph of the best fit line. Include three errors on the graph. Explain what those errors are and then explain how you would fix those errors to make the graph more accurate.
Answer #1

y = 0.245 x + 0.342

a)We have to substitute the absorbance values (y-axis value) in the above equation to get the concentration (in %) in form of x values.

0.342 = 0.245 x + 0.342

x = 0

3.861 = 0.245 x + 0.342

x = 14.36%

1.462 = 0.245 x + 0.342

x = 4.57%

b) To get the absorbance values, we have to  substitute the concentration values i.e 18% ,4% and 50% (x-axis value) in the above equation to get the absorbance values in form of y values.

y= 0.245 x + 0.342

y =0.245 * 0.18 + 0.342

= 0.39

y= 0.245 * 0.04 +0.342

= 0.35

y= 0.245 * 0.5 +0.342

= 0.46